Subject: APO 3.5m users committee minutes, 10/1/12
From: Suzanne Hawley
Submitted: Wed, 3 Oct 2012 20:40:38 -0700 (PDT)
Message number: 1290
(previous: 1289,
next: 1291
up: Index)
APO 3.5-m Users Committee Phonecon, 10/01/12
Attending: Suzanne Hawley, Mark Klaene, Al Harper, Cynthia Froning (for
John Bally), Scott Anderson, Remy Indebetouw, Jon Holtzman, Rene
Walterbos, Michael Strauss, Sean Moran, Bill Ketzeback, and Bruce
Gillespie.
Absent:
**********************************
User feedback and comments from institutional representatives:
o Chicago (Al Harper) - Al had nothing new to report.
o Washington (Scott Anderson) - Scott had nothing new to report.
o New Mexico State (Jon Holtzman) - Jon had nothing new to report.
o Princeton (Michael Strauss) - Michael had nothing new to report.
o Colorado (Cynthia Froning) - Cynthia had nothing new to report.
o Virginia (Remy Indebetouw) - Remy had nothing new to report.
o Johns Hopkins (Sean Moran) - Sean had nothing new to report.
**********************************
Discussion of telescope/instruments report:
Mark said that it has been relatively quiet since the last meeting, and
the telescope has been running well. The recent problems with vacuum
and cryo systems on DIS seem to be resolved, but we're keeping an eye
on the DIS slit viewer because of one or more episodes where its focus
seemed off. A recently installed ion pump for SPIcam failed, but we were
able to get a new one quickly and the instrument is back up and running.
We have been planning for an upgrade to the vacuum system for SPIcam
for later, but we've started it early with a new vacuum pump. There
seems to be some electrical pickup from the new pump, and more
electrical isolation is going to be added.
Mark Klaene's detailed report for the reporting period follows:
*************************
3.5-m Telescope and Instruments Highlights
09/04/12 through 09/28/12
0) Overview
Last week saw a large drop in our average dewpoint signaling an end to our
monsoon season. However a number of small disturbances continued to bring
some moisture into the area.
1) Telescope
The telescope performance has been nominal this period. Seeing conditions
did provide a couple of nights with 0.5" seeing at the echelle.
2) Instruments --
DIS has been operating nominally. Both red and blue cameras have been
functioning normally with good vacuum and temperature. There was one
instance of slit viewer image degradration. It is still under
investigation however we are convinced the slit viewer focus vs science
cameras offset is still correct.
Agile had no issues this period.
SPIcam's new ion pump failed causing the instrument to go offline for
about 9 days. No scheduled observing was lost. A new ion pump system was
purchased. This makes the ion pumps in DIS and SPIcam identical; we have
plans to implement identical controllers next year.
A new instrument cart was delivered for TripleSpec that improves mounting
and appears to fix the issue of the instrument coming off the kinematic
mounts. Final adjustments to the cart have been made and it is now in
routine use.
Echelle is operational. The old BL67 grating was measured by Newport
Optical. Observers had reported variations in blue flats and a problem
with the flip mirror motor driver was noticed. After a lot of reverse
engineering, we are close to designing a system around the current motor
controller to correct the problem. It is not yet confirmed this will fix
the variations in the flats. Note that both red and blue filter flats are
affected, although the high counts in the red make the variations less
noticeable.
NIC-FPS is operating in shared-risk observing mode due to occasional image
corruption issues.
**********************************
New echelle grating:
Jon Holtzman reported that the new grating is not showing the expected
improvement in throughput. The old grating was sent back to Newport for
comparison testing against the efficiency measurements taken there more
recently on the new grating. The old grating showed efficiencies somewhat
better than its original specifications, but not as high as the Newport
measurements for the new grating prior to its shipment to APO. Michael
Strauss asked if the new grating is to be considered the new baseline
setup, and Jon said yes, and this means that changes in calibrations will
be required to use it properly. We would like to hear feedback on this
new grating from echelle users while we're investigating the matter.
**********************************
Possible new imager options:
Suzanne, with help from Mark and Bill, described the status and plans for
obtaining a new visible imager for the telescope. GRIZLI, a dichroic
multi-band imager planned for development in collaboration with
Harvard-Smithsonian CfA, has stalled out -- we are now looking at an
alternative arrangement by working with Harvard alone, possibly employing
LSST chips through an collaboration with Brookhaven National Lab. A
preliminary optical design is being evaluated that yields a single 8- to
10-arcminute field of view, but there are issues involving pixel size
(small) versus the complexity of using a focal reducer to properly sample
the image. Suzanne said she would be interested in further ideas and
suggestions from the user community.
**********************************
Science input for BoG meeting:
Suzanne noted that we have received scientific highlights for the ARC
Board annual report from Chicago, Colorado, Virginia, and Princeton. We
hope that those from JHU, NMSU, and Washington will be forthcoming
quickly. Scott reported that UW slides will be sent today and Jon said
the same for NMSU.
**********************************
ATI proposal:
Suzanne, Mark, and Jon remarked that no one responded to the action item
from last month regarding looking for institutional participation in an
NSF ATI proposal to build a multi-fiber interface between the APOGEE
spectrograph and the 3.5-m telescope. A single fiber was recently run
between the APOGEE instrument and the NMSU 1-m telescope, and it has seen
first starlight. In essence, we would be able to feed the APOGEE
spectrograph with roughly 200 fibers collected at the 3.5-m during dark
time, either as IFUs or pluggable plates for individual target objects.
Suzanne stressed that developing this capability would bear significantly
on the future of the telescope and the ability to attract new consortium
partners. An idea for the "broader impact" part of the proposal that
could strengthen the chances of getting funding, is to offer a certain
amount of observing time on the telescope as "open" to U.S. graduate
students, especially in light of the recent recommendations of the NSF
Portfolio Review regarding future community access to mid-sized
telescopes, particularly for students.
The ATI proposal is due at the end of this month, so Suzanne needs
community and committee members' feedback immediately, especially on the
science case for the fiber feed to the 3.5m.
**********************************
Update on new partners:
Suzanne and Rene have previously asked for suggestions on which
institutions (and their contact persons) we should be approaching as
potential new ARC partners. We have received only a small amount of
feedback so far. Later this year, we will be placing a general
announcement of opportunity through the AAS newsletter, assuming that no
strong candidates are brought forward in the meantime.
**********************************
ACTION ITEMS:
[open from previous meeting]:
===> ACTION: Send new ARC partner suggestions to Suzanne and Rene (before
next meeting).
Status: Suzanne has received a small number of suggestions, and
is still looking for more. Please send your suggestions of an institution
and a contact person to her. Within two years Princeton will be leaving
ARC, and Chicago will also be looking at parting ways. This leaves 1/3 of
3.5-m time available and unpaid for. Therefore, we need to seriously
start soliciting new partners. A few feelers have produced some positive
feedback.
===> ACTION: Science slides for BOG meeting should be sent to Suzanne ASAP.
===> ACTION: NOW!! Users committee members should discuss APOGEE fiber
feed to 3.5m proposal with their users and forward feedback, and ideally
name(s) of interested users who would be willing to participate in an ATI
proposal to Suzanne, Mark and Fred Hearty (see email addresses on CC line
of this message).
**********************************
Next meeting: The next Users Committee phonecon will be on Monday, 5
November at 8:30 AM Pacific Time. The agenda and other materials will be
sent to the committee members during the preceding week.
APO APO APO APO APO Apache Point Observatory 3.5m APO APO APO
APO
APO This is message 1290 in the apo35-general archive. You can find
APO the archive on http://www.astro.princeton.edu/APO/apo35-general/INDEX.html
APO To join/leave the list, send mail to apo35-request@astro.princeton.edu
APO To post a message, mail it to apo35-general@astro.princeton.edu
APO
APO APO APO APO APO APO APO APO APO APO APO APO APO APO APO APO